Exeter has a material standard of 1 pound per unit of output. Each pound has a standard price of $27 per pound. During July, Exeter paid $137,000 for 4,990 pounds, which it used to produce 4,760 units. What is the direct materials price variance? (Do not round your intermediate calculations.)

$8,480 unfavorable

$1,370 favorable

$9,850 unfavorable

$2,270 unfavorable
Answer #1

Materials Price Variance = (SP-AP)*AQ

= (27 - 137,000/4990)*4990

= (27 - 27.45) * 4990

= 2270 Unfavorable

Option D is the answer